# Multi-Excitation Projective Simulation with a Many-Body Physics Inspired Inductive Bias

### v2

@@ VI First Steps Towards Quantum mePS
As we described in Section IV.1, the dynamics of such quantum systems are described by time evolution operators of the form $e^{itH}$ (or more generally, $\mathcal{T}e^{\int_{t_{1}}^{t_{2}}H(t)\mathrm{d}t}$, where $\mathcal{T}$ is the time-ordering operator). The coefficients $h(\{j_{1},\ldots,j_{i}\},\{k_{1},\ldots,k_{o}\})\in\mathbb{C}$ in Eq. (4) are then the trainable parameters or functions of the trainable parameters (such as tunable tunnelling amplitudes and 2-body interaction couplings).
@@
One issue to consider is that physical Hamiltonians $H$ are Hermitian, i.e. $H^{\dagger}=H$. This has the consequence that $h(\{j_{1},\ldots,j_{i}\},\{k_{1},\ldots,k_{o}\})=\overline{h(\{j_{1},\ldots,j_{i}\},\{k_{1},\ldots,k_{o}\})}$, meaning that the transition amplitude to go forward is just as large as the amplitude to go backward. This issue is already present in the quantization of basic PS and was addressed in  [14] by using dissipation (or other irreversible, open quantum system evolutions) to obtain a broader class of parametrized time evolutions.
Another approach is to introduce an extra semi-classical degree of freedom that breaks the symmetry by acting as a clock. One inspiration for how to do this comes from integrated interferometer chips. Here, the photons have a definite velocity in the lateral direction of the photonics chip, but perform quantum walks in the transversal direction [48]. This approach was applied in a recent proposal for a quantum PS with a single photon [47].
It was mentioned in Remark 8 that our choice in classical mePS about two excitations in the same clip does not faithfully dequantize the behavior of any quantum particles. It would be interesting to analyze how the decision-making process is influenced by also being physically faithful in this regard. For example, the phenomenon of Pauli pressure in fermions could prevent an already excited clip from being excited again, potentially putting the second excitation to better use somewhere else. This has the consequence that the random walks of fermions depend on each other, even if the Hamiltonian has no interaction terms.
@@ VII Discussion and Outlook
In this paper, we introduced an XAI method called mePS, which allows us to model chains of thoughts as random walks of several particles on a hypergraph. The use of several particles allows for the representation of thoughts relying on the combination of several elementary concepts simultaneously, revealing and exploiting the composite structure of thoughts and thus greatly improving model interpretability. . This added flexibility is a stepping stone in developing systematic methods that let us model domain knowledge via the structure of the hypergraph and attach concepts to clusters of relevant vertices on the hypergraph. A new definition for dynamic hypergraph involving foliations was also introduced to model the agent’s training history and it was suggested that mePS could serve as a hypergraph generator with applications to hypergraph visualization.
To reduce the exponential complexity of a naive implementation of mePS to a low-degree polynomial complexity, we defined an inductive bias. This inductive bias is a classical analogue of the time evolution of quantum many-body systems. This inductive bias includes a cutoff regarding how many particles can participate in an interaction. We proved that our inductive bias indeed leads to a polynomial complexity, with the degree given by the interaction cutoff. We believe that our inductive bias does not severely restrict the potential of mePS agents in many scenarios. This belief is motivated by the fact that humans can also only combine a handful of concepts simultaneously. Nonetheless, humans display an unmatched ability to quickly adapt to a wide range of environments. Furthermore, this focus on only a few concepts at a time could be seen as a conceptual analogue of the attention mechanism used in LLMs [52, 53], at least for the case of sparse attention matrices. Similarly to human attention, the LLM attention mechanism is known to perform very well [54].
@@
Quantum many-body systems have additional properties which we did not consider in this work. One important such property is that particles can only directly interact if they are physically close. Particles classified as fermions tend to avoid each other, and the related mechanisms are known under names such as Pauli-exclusion, Pauli-pressure, and excluded volume. Nonetheless, local elementary interactions of few particles give rise to most phenomena known in physics. Therefore, it might be worthwhile to model an additional inductive bias which formalizes a notion of distance between clips, and restricts $h^{(n)}_{(i,o)}$ to only be non-zero for close excitations.
@@
Another important property of quantum many-body systems concerns the question of what happens if one tries to put an excitation on an already excited clip. As explained in Remark 8, we made an unphysical choice motivated by the interpretation of clips as concepts. Therefore, it would be interesting to investigate what would happen if one instead mimicked physics in this aspect. For fermions, putting an excitation in an already excited clip would result in an empty clip. For bosons, several excitations could be in the same clip, meaning one would first have to model an extension of mePS that allows several excitations on the same clip.
@@
The binary nature of our excitation configurations suggests the existence of potential relations to the field of Neuro-Symbolic (NeSy) logic [56, 57]. Indeed, one can read the presence of an excitation on a clip as a truth state that the concept represented by the clip is currently relevant or applicable. Investigating these relations might lead to a fruitful cross-fertilization between the two fields. However, it is also important to emphasize the differences. An important ambition of our mePS scheme is to model various chain-of-thought processes, not just those relying on formal logic to process facts of the environment. These also include thought processes that underly irrational or bounded rational decisions, as studied in psychology and the decision sciences. Furthermore, mePS comes with a natural update rule based on h-values that could also be applied to NeSy.
@@
On the numerical side, compiling the for-loops for or partially parallelizing the deliberation process, or using advanced Monte Carlo Simulation [58, 59] software might significantly speed up mePS agents. For Python implementations, a first route towards this goal might be to use Cython [60] or just-in-time compiling modules like Numba [61]. To numerically profit from our inductive biases, it is important to sample transitions in a way that does not require iterating over the full power set of clips. We already formulated one method for sampling transitions, but we have no guarantee that it is the best possible implementation. Indeed, we appealed very little to results from the mathematical literature on hypergraphs such as hypergraph expansion techniques (star, cluster, line, etc.), Laplacian spectral clustering techniques, factorization of hypergraph matrix representations, and other hierarchical partitionings of hypergraphs in developing mePS  [22, 23, 24]. We also assumed the clips within each hyperedge were of equal importance, which may not be the case in situations where some information or properties of the data are privileged over others, so incorporating clip weights into the training process could be beneficial in this regard  [24]. Further exploration into the hypergraph literature will likely yield many improvements for our mePS methodology and should be considered as an important next step in the development of mePS; chiefly for the scalability of the model.
@@
Furthermore, we employed a dual reward mechanism to avoid the credit assignment problem but we did not explore other potential mitigation strategies that could alleviate this problem. A search for systematic methods to find good initialization strategies, or to adapt Imitation Learning methods to our setting [62, 63] would prove useful.
@@ Appendix A Additional Computer Maintenance Training Details
Appendix B The Relation Between Hypergraphs
@@
$\sum_{\begin{subarray}{c}(i,o)\in IO,\ (C_{\mathrm{in}}^{(i)}\rightarrow C_{\mathrm{out}}^{(o)})\in E^{(i,o)}\\ \text{such that }C_{\mathrm{in}}^{(i)}\subset C_{\mathrm{in}},\ C_{\mathrm{out}}=(C_{\mathrm{in}}\setminus C_{\mathrm{in}}^{(i)})\cup C_{\mathrm{out}}^{(o)}\end{subarray}}h_{(i,o)}(C_{\mathrm{in}}^{(i)},C_{\mathrm{out}}^{(o)})$ (15)
@@
$\sum_{\begin{subarray}{c}(i,o)\in IO,\ (C_{\mathrm{in}}^{(i)}\rightarrow C_{\mathrm{out}}^{(o)})\in E^{(i,o)}\\ \text{such that }C_{\mathrm{in}}^{(i)}\subset C_{\mathrm{in}},\ C_{\mathrm{out}}=(C_{\mathrm{in}}\setminus C_{\mathrm{in}}^{(i)})\cup C_{\mathrm{out}}^{(o)}\end{subarray}}p(C_{\mathrm{in}}^{(i)}\rightarrow C_{\mathrm{out}}^{(o)})$ (19)
@@
$\sum_{\begin{subarray}{c}(i,o)\in IO,\ (C_{\mathrm{in}}^{(i)}\rightarrow C_{\mathrm{out}}^{(o)})\in E^{(i,o)}\\ \text{such that }C_{\mathrm{in}}^{(i)}\subset C_{\mathrm{in}},\ C_{\mathrm{out}}=(C_{\mathrm{in}}\setminus C_{\mathrm{in}}^{(i)})\cup C_{\mathrm{out}}^{(o)}\end{subarray}}\frac{h_{(i,o)}(C_{\mathrm{in}}^{(i)},C_{\mathrm{out}}^{(o)})}{\mathcal{N}}$ (27)
@@
$\sum_{\begin{subarray}{c}(i^{\prime},o^{\prime})\in IO,\ (D_{\mathrm{in}}^{(i^{\prime})}\rightarrow D_{\mathrm{out}}^{(o^{\prime})})\in E^{(i^{\prime},o^{\prime})}\\ \text{such that }D_{\mathrm{in}}^{(i^{\prime})}\subset C_{\mathrm{in}},\ C_{\mathrm{out}}=(C_{\mathrm{in}}\setminus D_{\mathrm{in}}^{(i^{\prime})})\cup D_{\mathrm{out}}^{(o^{\prime})}\end{subarray}}h_{(i^{\prime},o^{\prime})}(D_{\mathrm{in}}^{(i^{\prime})},D_{\mathrm{out}}^{(o^{\prime})})$ (30)
@@
We point out that the proof can be adapted to the Inductive Biases 2:
@@
Appendix C Further complexity results and proofs
In this appendix, we provide additional complexity estimates and proofs.
First, we provide a simple scenario for Inductive Bias 1 that contains cycles, and therefore allows for infinitely long random walks.

### v3

@@ VI.2 First Steps Towards Quantum mePS
As we described in Section VI.1, the dynamics of such quantum systems are described by time evolution operators of the form $e^{-itH}$ (or more generally, $\mathcal{T}e^{-i\int_{t_{1}}^{t_{2}}H(t)\mathrm{d}t}$, where $\mathcal{T}$ is the time-ordering operator).
We emphasize that most of the considerations in Section VI.1 served to formally develop a classical and discrete analogue. On quantum simulation hardware, we can directly define a parametrized Hamiltonian $H$ of the form in Eq. (8) and define the deliberation as its time evolution, which is a continuous time quantum walk.
The coefficients $h(\{j_{1},\ldots,j_{i}\},\{k_{1},\ldots,k_{o}\})\in\mathbb{C}$ in Eq. (8) are then the trainable parameters or functions of the trainable parameters (such as tunable tunnelling amplitudes and 2-body interaction couplings).
@@
One issue to consider is that physical Hamiltonians $H$ are Hermitian, i.e. $H^{\dagger}=H$. This has the consequence that $h(\{j_{1},\ldots,j_{i}\},\{k_{1},\ldots,k_{o}\})=\overline{h(\{k_{1},\ldots,k_{o}\},\{j_{1},\ldots,j_{i}\},)}$, meaning that the transition amplitude to go forward is just as large as the amplitude to go backward. This issue is already present in the quantization of basic PS and was addressed in  [14] by using dissipation (or other irreversible, open quantum system evolutions) to obtain a broader class of parametrized time evolutions.
Another approach is to introduce an extra semi-classical degree of freedom that breaks the symmetry by acting as a clock. One inspiration for how to do this comes from integrated interferometer chips. Here, the photons have a definite velocity in the lateral direction of the photonics chip, but perform quantum walks in the transversal direction [47]. This approach was applied in a recent proposal for a quantum PS with photons [46].
It was mentioned in Remark 8 that our choice in classical mePS about two excitations in the same atomic clip does not faithfully dequantize the behavior of any quantum particles. It would be interesting to analyze how the decision-making process is influenced by also being physically faithful in this regard. For example, the phenomenon of Pauli pressure in fermions could prevent an already excited atomic clip from being excited again, potentially putting the second excitation to better use somewhere else. This has the consequence that the random walks of fermions depend on each other, even if the Hamiltonian has no interaction terms.
@@ VII Discussion and Outlook
In this paper, we introduced an XAI framework called mePS, which allows us to model chains of thoughts as random walks of several particles on a hypergraph. The use of several particles allows for the representation of thoughts relying on the combination of several elementary concepts simultaneously, revealing and exploiting the composite structure of thoughts and thus greatly improving model interpretability. This added flexibility is a stepping stone in developing systematic methods that let us model domain knowledge via the structure of the hypergraph and attach concepts to clusters of relevant vertices on the hypergraph. A new definition for dynamic hypergraph was also introduced to model the agent’s training history and serve as a tool for explainability by connecting with the hypergraph visualization literature.
To reduce the exponential complexity of a naive implementation of mePS to a low-degree polynomial complexity, we defined an inductive bias. This inductive bias is a classical analogue of the time evolution of quantum many-body systems. This inductive bias includes a cutoff regarding how many particles can participate in an interaction. We proved that our inductive bias indeed leads to a polynomial complexity, with the degree given by the interaction cutoff. We believe that our inductive bias does not severely restrict the potential of mePS agents in many scenarios. This belief is motivated by the fact that humans can also only combine a handful of concepts simultaneously. Nonetheless, humans display an unmatched ability to quickly adapt to a wide range of environments.
@@
Quantum many-body systems have additional properties which we did not consider in this work. One important such property is that particles can only directly interact if they are physically close. Particles classified as fermions tend to avoid each other, and the related mechanisms are known under names such as Pauli-exclusion, Pauli-pressure, and excluded volume. Nonetheless, local elementary interactions of few particles give rise to most phenomena known in physics. Therefore, it might be worthwhile to model an additional inductive bias which formalizes a notion of distance between atomic clips, and restricts $h^{(n)}_{(i,o)}$ to only be non-zero for close excitations.
@@
Another important property of quantum many-body systems concerns the question of what happens if one tries to put an excitation on an already excited atomic clip. As explained in Remark 8, we made an unphysical choice motivated by the interpretation of atomic clips as concepts. Therefore, it would be interesting to investigate what would happen if one instead mimicked physics in this aspect. For fermions, putting an excitation in an already excited atomic clip would result in an empty atomic clip. For bosons, several excitations could be in the same atomic clip, meaning one would first have to model an extension of mePS that allows several excitations on the same atomic clip.
@@
The binary nature of our excitation configurations suggests the existence of potential relations to the field of Neuro-Symbolic (NeSy) logic [52, 53]. Indeed, one can read the presence of an excitation on an atomic clip as a truth state that the concept represented by the atomic clip is currently relevant or applicable. Investigating these relations might lead to a fruitful cross-fertilization between the two fields. However, it is also important to emphasize the differences. An important ambition of our mePS scheme is to model various chain-of-thought processes, not just those relying on formal logic to process facts of the environment. These also include thought processes that underly irrational or bounded rational decisions, as studied in psychology and the decision sciences. Furthermore, mePS comes with a natural update rule based on h-values that could also be applied to NeSy.
@@
On the numerical side, compiling the for-loops for or partially parallelizing the deliberation process, or using advanced Monte Carlo Simulation [54, 55] software might significantly speed up mePS agents. For Python implementations, a first route towards this goal might be to use Cython [56] or just-in-time compiling modules like Numba [57]. To numerically profit from our inductive biases, it is important to sample transitions in a way that does not require iterating over the full power set of atomic clips. We already formulated one method for sampling transitions, but we have no guarantee that it is the best possible implementation. Indeed, we appealed very little to results from the mathematical literature on hypergraphs such as hypergraph expansion techniques (star, cluster, line, etc.), Laplacian spectral clustering techniques, factorization of hypergraph matrix representations, and other hierarchical partitionings of hypergraphs in developing mePS  [22, 23, 24]. We also assumed the atomic clips within each hyperedge were of equal importance, which may not be the case in situations where some information or properties of the data are privileged over others, so incorporating atomic clip weights into the training process could be beneficial in this regard  [24]. Further exploration into the hypergraph literature will likely yield many improvements for our mePS methodology and should be considered as an important next step in the development of mePS; chiefly for the scalability of the model.
@@
Furthermore, we employed a dual reward mechanism to avoid the credit assignment problem but we did not explore other potential mitigation strategies that could alleviate this problem. A search for systematic methods to find good initialization strategies, or to adapt Imitation Learning methods to our setting [58, 59] might prove useful.
@@ Appendix A Additional Computer Maintenance Training Details
Appendix B Modifications of the Inductive Bias
In the main text, we mentioned two modifications of Inductive Bias 2FF which guarantee low-order polynomial random walk lengths in the depth and width of the layered ECM. Now, we formulate their precise definition:
Inductive Bias 2SF and 2DP.
For weighted, layered feed-forward many-body hypergraphs with layers $(L_{1},\dots,L_{D})$, we introduce the following two modifications of Inductive Bias 1 and 2FF:
  1. SF.
    Let $\ell:V\rightarrow\{1,2,\dots D\}$ be the function that maps each atomic clip to the layer it is in. Then, given an excitation configuration $\{c_{m_{1}},\dots,c_{m_{x}}\}$, with the labelling such that the layers satisfy $\ell(c_{m_{1}})=\cdots=\ell(c_{m_{n}})<\ell(c_{m_{n+1}})\leq\cdots\leq\ell(c_{m_{x}})$, the relevant $h_{(i,o)}(C_{\mathrm{in}},C_{\mathrm{out}})$-values for this configuration are restricted to those with $C_{\mathrm{in}}\subset\{c_{m_{1}},\dots,c_{m_{n}}\}$.
  2. DP.
In Appendix C, we will show that all inductive biases, including 2SF and 2DP, satisfy Definitions 2 and 5 when the standard probability assignment is used.
In Appendix D, we will prove the promised upper bounds.
Now, to demonstrate how these inductive biases work, we revisit Example 11, and this time include Inductive Biases 2SF and 2DP.
Figure 9: An example illustrating random walk steps under different inductive biases, compare with Example 11. Excited atomic clips are shown in red. The sampled hyperedge is shown in blue. Subfigure a) shows a deliberation step which is only allowed under Inductive Bias 1, because its codomain is in two layers. Also, it shows that an excitation moving into an occupied atomic clip gets discarded. Subfigure b) shows a typical transition under Inductive Biases 2FF and 2SF. Notice that it can lead to excitations being spread over several layers. Subfigure c) shows a typical transition under Inductive Bias 2DP. It discards uninvolved excitations, and therefore only contains excitations in the codomain of the hyperedge.
Example 16 (Example 11 including Inductive Biases 2SF and 2DP).
Consider a simple 2-layer setting, with 4 atomic clips in each layer, see Figure 9: $V=L_{1}\cup L_{2}$, with $L_{1}=\{c_{1},c_{2},c_{3},c_{4}\}$ and $L_{2}=\{c^{\prime}_{1},c^{\prime}_{2},c^{\prime}_{3},c^{\prime}_{4}\}$. We only consider h-values with the same number of incoming and outgoing excitation numbers, and let no more than two excitations interact. That means $IO=\{(1,1),(2,2)\}$. Our current excitation configuration is $\{c_{1},c_{2},c_{3}\}$, meaning that we currently have an excitation in each of the atomic clips $c_{1}$, $c_{2}$, and $c_{3}$.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    $h_{(1,1)}(c_{m},c^{\prime}_{j})$ such that $j\in\{1,2,3,4\}$ and $m\in\{1,2,3\}$
  5. 5.
This list gets turned into probabilities, in our example by applying the softmax-function to the full list. Say, we sample $h_{(2,2)}(\{c_{2},c_{3}\},\{c_{1},c^{\prime}_{1}\})$ and apply it to our current configuration $\{c_{1},c_{2},c_{3}\}$. First, we remove the excitations in $c_{2}$ and $c_{3}$, giving us the configuration $\{c_{1}\}$. Next, we put excitations into $c_{1}$ and $c^{\prime}_{1}$. However, $c_{1}$ already carries an excitation. We just keep this excitation as it is. So our next excitation configuration is $\{c_{1},c^{\prime}_{1}\}$. Note that our rule for dealing with already occupied atomic clips led to a reduction in the total number of excitations.
Our layered Inductive Biases 2FF and 2SF differ from the previous situation in that the relevant many-body h-values are only items 1 and 4 from the numbered list above. Now, say that we sampled $h_{(2,2)}(\{c_{1},c_{2}\},\{c^{\prime}_{2},c^{\prime}_{3}\})$ and apply it to our current configuration $\{c_{1},c_{2},c_{3}\}$. First, we remove the excitations in $c_{1}$ and $c_{2}$, giving us the configuration $\{c_{3}\}$. Next, we insert excitations in $c^{\prime}_{2},c^{\prime}_{3}$, giving us the full next excitation configuration $\{c^{\prime}_{2},c^{\prime}_{3},c_{3}\}$. We observe that while the feed-forward condition forces all excitations that move to move one layer forward, it allows excitations to stay behind in their old atomic clip in the old layer. Consider now an additional layer $L_{3}$. While Inductive Bias 2FF allows us to continue with any transition $C_{\mathrm{in}}\rightarrow C_{\mathrm{out}}$ that has $C_{\mathrm{in}}\subset\{c^{\prime}_{2},c^{\prime}_{3}\}$ or $C_{\mathrm{in}}=\{c_{3}\}$, Inductive Bias 2SF forces us to remove $c_{3}$ first. For Inductive Bias 2SF, the relevant many-body h-values are therefore just those of the form $h_{(1,1)}(c_{3},c^{\prime}_{j})$ with $j=1,\dots 4$.
Now, consider Inductive Bias 2DP acting on $\{c_{1},c_{2},c_{3}\}$. While it has the same list of relevant $h$-values as Inductive Bias 2FF, it applies the transitions another way. Again, assume that we sampled $h_{(2,2)}(\{c_{1},c_{2}\},\{c^{\prime}_{2},c^{\prime}_{3}\})$ and apply it to $\{c_{1},c_{2},c_{3}\}$. Again, we remove the ingoing excitations $c_{1}$ and $c_{2}$, giving us $\{c_{3}\}$. Next, we insert excitations in $c^{\prime}_{2}$ and $c^{\prime}_{3}$, giving us $\{c^{\prime}_{2},c^{\prime}_{3},c_{3}\}$. Furthermore, $c_{3}$ is neither an ingoing nor an outgoing atomic clip of $h_{(2,2)}(\{c_{1},c_{2}\},\{c^{\prime}_{2},c^{\prime}_{3}\})$, so we discard its excitation. This gives us the full next excitation configuration $\{c^{\prime}_{2},c^{\prime}_{3}\}$. As we see, Inductive Bias 2DP enforces that after a transition, all excitations are found in the same layer.
When modeling agents with Inductive Bias 2SF, there is one important consequence to keep in mind: In human decision-making, a common theme is to write down some intermediate results, and only use them much later when they are deemed relevant. An example would be the derivation of several independent lemmas, all of which get used in proving a theorem. Since in Inductive Bias 2SF the shallowest excitations are removed first, one should introduce copies of their atomic clips in deeper layers to not lose the knowledge/concepts they represent in later steps.
Appendix C The Relation Between Hypergraphs
@@
$\sum_{\begin{subarray}{c}(i,o)\in IO,\ (C_{\mathrm{in}}^{(i)}\rightarrow C_{\mathrm{out}}^{(o)})\in E^{(i,o)}\\ \text{such that }C_{\mathrm{in}}^{(i)}\subset C_{\mathrm{in}},\ C_{\mathrm{out}}=(C_{\mathrm{in}}\setminus C_{\mathrm{in}}^{(i)})\cup C_{\mathrm{out}}^{(o)}\end{subarray}}h_{(i,o)}(C_{\mathrm{in}}^{(i)},C_{\mathrm{out}}^{(o)})$
@@
$\sum_{\begin{subarray}{c}(i,o)\in IO,\ (C_{\mathrm{in}}^{(i)}\rightarrow C_{\mathrm{out}}^{(o)})\in E^{(i,o)}\\ \text{such that }C_{\mathrm{in}}^{(i)}\subset C_{\mathrm{in}},\ C_{\mathrm{out}}=(C_{\mathrm{in}}\setminus C_{\mathrm{in}}^{(i)})\cup C_{\mathrm{out}}^{(o)}\end{subarray}}p(C_{\mathrm{in}}^{(i)}\rightarrow C_{\mathrm{out}}^{(o)})$
@@
$\sum_{\begin{subarray}{c}(i,o)\in IO,\ (C_{\mathrm{in}}^{(i)}\rightarrow C_{\mathrm{out}}^{(o)})\in E^{(i,o)}\\ \text{such that }C_{\mathrm{in}}^{(i)}\subset C_{\mathrm{in}},\ C_{\mathrm{out}}=(C_{\mathrm{in}}\setminus C_{\mathrm{in}}^{(i)})\cup C_{\mathrm{out}}^{(o)}\end{subarray}}\frac{h_{(i,o)}(C_{\mathrm{in}}^{(i)},C_{\mathrm{out}}^{(o)})}{\mathcal{N}}$
@@
$\sum_{\begin{subarray}{c}(i^{\prime},o^{\prime})\in IO,\ (D_{\mathrm{in}}^{(i^{\prime})}\rightarrow D_{\mathrm{out}}^{(o^{\prime})})\in E^{(i^{\prime},o^{\prime})}\\ \text{such that }D_{\mathrm{in}}^{(i^{\prime})}\subset C_{\mathrm{in}},\ C_{\mathrm{out}}=(C_{\mathrm{in}}\setminus D_{\mathrm{in}}^{(i^{\prime})})\cup D_{\mathrm{out}}^{(o^{\prime})}\end{subarray}}h_{(i^{\prime},o^{\prime})}(D_{\mathrm{in}}^{(i^{\prime})},D_{\mathrm{out}}^{(o^{\prime})})$
@@
We point out that the proof can be adapted to the Inductive Biases 2FF, 2SF and 2DP:
@@
Appendix D Detailed Complexity Results and Proofs
So far, our resource estimates do not distinguish between our inductive biases. However, we already hinted at the fact that they have crucial complexity differences with regard to the maximal deliberation time. To see the difference, we derive bounds on the deliberation time.
At first, we point out that Inductive Bias 1 still allows for deliberations to be arbitrarily long: For any attainable excitation configuration $\{c_{m_{1}},\dots,c_{m_{x}}\}$, if it is possible to combine transitions relevant for this configuration to a cycle leading again to $\{c_{m_{1}},\dots,c_{m_{x}}\}$, then this cycle can lead to arbitrarily long deliberation times. A simple class of Inductive Bias 1 agents that have such cycles is presented in the following proposition: